OpenStudy (anonymous):

Can anyone help me with this question give exact and approximate solutions to three decimal places. x^2-5x+5=0 The exact solutions are x= the approximate solutions to 3 decimal places are x=

jhonyy9 (jhonyy9):

- if x,y are solutions - x=(5+(25-20)*(1/2))/2=((5+5*(1/2))/2=(5+2,41)/2=7,41/2= =3,705 y=((5-(25-20)*(1/2))/2=(5-2,41)/2=2,59/2=1,295